# MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@ Abstract
Interest in quantum machine learning is increasingly growing due to the possibility of developing efficient solutions to problems that are difficult to tackle with classical methods. In this context, the research work presented here focuses on the use of quantum machine learning techniques for image classification tasks. We exploit a quantum extreme learning machine by taking advantage of its rich feature map provided by the quantum reservoir substrate. We systematically analyse different phases of the quantum extreme learning machine process, from the dataset preparation to the image final classification. In particular, we investigate the impact of encoding through a Principal Component Analysis and the use of Auto-Encoders, as well as the dynamics of the model through the use of different Hamiltonians for the quantum reservoir. Our results show that the introduction of a quantum reservoir systematically improves the accuracy of the classifier. Additionally, while different encodings can lead to significantly different performances, Hamiltonians with varying degrees of connectivity exhibit the same discrimination rate, provided they are interacting.
@@ 1   Introduction
Since the dawn of the new millennium, quantum computing has steadily gained prominence, emerging as a groundbreaking technology with the potential to revolutionize various sectors, from chemistry and cryptography to finance and artificial intelligence.
@@
One particularly exciting and rapidly evolving field within quantum computing is quantum machine learning. This innovative domain, which lies at the intersection of quantum computing and machine learning, aims to harness quantum algorithms and data structures to tackle complex problems and achieve breakthroughs that surpass the capabilities of classical machine learning methods.
In this context, this paper explores the use of a Quantum Extreme Learning Machine (QELM) [1, 2, 3, 4, 5], leveraging the advanced feature map provided by the quantum reservoir substrate. QELMs demonstrate substantial potential by harnessing the complex dynamics of a quantum system within an exponentially large Hilbert space. A further advantage of QELMs is that only the final (output) layer of the model requires training, significantly reducing computational resource requirements. This results in a faster and more cost-effective training process compared to models that necessitate full training.
@@
Throughout this paper, we study the architecture and the performance of the QELM for image classification tasks. Due to the limitations arising from the reduced number of qubits available today and the small decoherence times of current noisy quantum devices [6], typical datasets, characterised by a large number of features, cannot be directly fed to the quantum reservoirs and must be necessarily compressed into a lower dimensional representation. In the context of QELM, this is usually achieved by linear reduction algorithms, such as Principal Component Analysis (PCA). In this work, we perform a step forward and exploit the more complex representation power of non-linear neural network architectures such as the Auto-Encoders.
@@
Subsequently, since it is necessary to convert classical information into quantum states that the quantum system can process, we explore various encoding methods. Finally, we compare the results obtained from different Hamiltonians generating the natural dynamics of the reservoir.
@@
The paper is structured as follows. Section 2 provides a description of the model used. Sections 3 and 4 cover techniques for reducing data complexity and methods for encoding classical input information, respectively. In Section 5, we present three different Hamiltonian models. Section 6 details the measurement process and the classical machine learning classifier. Finally, Section 7 discusses the model’s performance and results, followed by the conclusion.
@@ 2   Overview of Quantum Extreme Learning Machines
A schematic representation of the model is shown in Fig.1 and its workflow can be summarized as follows:
  • Feature reduction: the original data is compressed into a lower dimensional latent representation to match the limited number of available qubits of the reservoir. This can be achieved using either linear Principal Component Analysis (PCA) or a nonlinear Auto-Encoder (AE).
@@
Figure 1: A schematic representation of the QELM.
@@ 3   Techniques for reducing data complexity
In depth, we examine two distinct approaches for reducing the number of dimensions of the features: the Principal Component Analysis (PCA) and the Auto-Encoder (AE) [9, 10, 11, 12].
@@
In particular, we used two different algorithms of auto-encoders that we name $AE_{1}$ and $AE_{2}$. Both utilize in the encoder convolutional and max-pooling layers for feature extraction, while convolutional and up-sampling layers are used in the decoder for the reconstruction of the input data. A fully-connected layer maps the aforementioned features into a latent-space representation which is then fed to the quantum algorithm. The fully-connected layer employs a sigmoid activation function (the outputs of which are real values between 0 and 1) while the convolutional layers rely on the rectified linear unit. The two auto-encoders mainly differ in their expressivity potential, with $AE_{1}$ characterised by a number of trainable parameters larger than the one in $AE_{2}$. More details on the architectures of $AE_{1}$ and $AE_{2}$ can be found in Appendix A.
@@ 4   Encoding strategies of classical data into the QELM
|$\ket{\vec{x}}=\bigotimes_{i=1}^{\left\lceil N/2\right\rceil}\left(\cos(\frac{x_{2i-1}}{2})\ket{0}+e^{ix_{2i}}\sin(\frac{x_{2i-1}}{2})\ket{1}\right)$ (2)
where $x_{i}$ represents the elements of the feature vector $\vec{x}=[x_{1},...,x_{N}]^{T}\in\mathbb{R}^{N}$ and $N$ is the number of qubits. We chose to normalize each feature in the interval $[0,\pi]$, thus effectively covering, for each qubit, half of the corresponding Bloch sphere [3], with the angles $\theta$ and $\phi$ identified with pairs of features ($x_{2i-1}$, $x_{2i}$).
@@
A simpler case is represented by a single qubit encoding which maps each feature into one qubit, without exploiting the relative phase, namely with $\phi=0$. In this case, the initial state is prepared according to
|$\ket{\vec{x}}=\bigotimes_{i=1}^{N}\left(\cos(\frac{x_{i}}{2})\ket{0}+\sin(\frac{x_{i}}{2})\ket{1}\right).$ (3)
@@
|$\ket{\vec{x}}=\bigotimes_{i=1}^{\left\lceil N/2\right\rceil}\left(\sqrt{x_{2i-1}}\ket{0}+e^{ix_{2i}}\sqrt{1-x_{2i-1}})\ket{1}\right)$ (4)
@@
|$\ket{\vec{x}}=\bigotimes_{i=1}^{\left\lceil N/2\right\rceil}\frac{1}{\sqrt{x_{2i-1}^{2}+x_{2i}^{2}}}\left(x_{2i-1}\ket{0}+x_{2i}\ket{1}\right).$ (5)
@@ 5   Hamiltonian models describing the quantum layer
In this section, we explore the influence of three distinct Hamiltonians, which govern the time evolution of the initial state containing the encoded features, on the classification accuracy. In particular, we consider the time-dependent Hamiltonian studied in [3], and two time-independent operators, one with the qubit lying on a fully connected graph, with all-to-all interactions [19] and another with qubit on a line, with only nearest neighbor interactions [20].
@@
Here $\sigma_{x}^{(i)}$, $\sigma_{z}^{(i)}$, and $\sigma_{z}^{(j)}$ are the Pauli operators acting on the site of the $i$,$j$-th qubit, while $J_{1}^{i,j}=\frac{J_{0}}{|i-j|^{\alpha}}$ represents the long-range interaction between the $i$-th and $j$-th qubits. As in [3], the scale of the interaction is set by $J_{0}=0.06$ and is damped for far-away qubits, with $\alpha=1.51$ controlling the rate of damping. $B_{1}=3.05$ and the time evolution is performed using a $\Delta t=50T$ ($T=2T_{1}=1$, setting the units for both time and inverse energy).
In the second case, we used the Hamiltonian $H_{2}$ [19]
@@
Here the coefficients $J_{2}^{i,j}$ and $B_{2}^{(i)}$ are sampled from the Gaussian distributions $N(0.75,0.1)$ and $N(1,0.1)$, respectively. The time evolution is implemented with a time $\Delta t=20$. [19]
In the third case, the following Hamiltonian $H_{3}$ [20] was used
@@
where we set the parameters in the chaotic regime: $J_{3}=-1$, $B_{3x}=0.7$ and $B_{3z}=1.5$. Also for $H_{3}$ the time evolution is performed using a $\Delta t=20$.[20]
Using the Hamiltonians defined above, we implemented the time evolution of the quantum system using the QuTiP package [21, 22].
@@ 7.1   PCA vs Autoencoders
As naively expected, the simpler linear PCA shows a lower accuracy compared to the neural-network-based algorithms, $AE_{1}$ and $AE_{2}$, and $AE_{1}$ performs better than $AE_{2}$ due to its higher capacity. We notice that, after a fast initial growth, the accuracy starts to saturate around $N=9$ qubits to $\sim 98\%$ for the test set and slightly higher for the training set. As the number of qubits increases, the differences among the classification performances of the three algorithms get reduced. This behaviour can be ascribed to the corresponding increase in the dimension of the latent representation, combined with the relative simplicity of the MNIST dataset. As such, we expect the gap between the autoencoders and the PCA to persist even for large $N$ for more complex datasets.
Figure 2: Training (left panel) and testing (right panel) accuracy as a function of the number of qubits for different feature reduction schemes: PCA (blue) and two autoencoders (green and red curves). The time evolution and the encoding have been performed with the Hamiltonian $H_{1}$ and the dense angle, respectively.

### v3

@@ Abstract
Interest in quantum machine learning is increasingly growing due to its potential to offer more efficient solutions for problems that are difficult to tackle with classical methods. In this context, the research work presented here focuses on the use of quantum machine learning techniques for image classification tasks. We exploit a quantum extreme learning machine by taking advantage of its rich feature map provided by the quantum reservoir substrate. We systematically analyze different phases of the quantum extreme learning machine process, from the dataset preparation to the image final classification. In particular, we have tested different encodings, together with Principal Component Analysis, the use of Auto-Encoders, as well as the dynamics of the model through the use of different Hamiltonians for the quantum reservoir. Our results show that the introduction of a quantum reservoir systematically improves the accuracy of the classifier. Additionally, our findings indicate that variations in encoding methods can significantly influence performance and that Hamiltonians with distinct structures exhibit the same discrimination rate, depending on how their eigenstates are related to the encoding and measurement basis.
@@ 1   Introduction
Since the early 2000s, quantum computing has steadily gained prominence, emerging as a groundbreaking technology with the potential to revolutionize various sectors, including chemistry, cryptography, finance, and artificial intelligence [1, 2, 3, 4]. In recent years, significant technological advancements [5, 6, 7, 8] have accelerated interest in this field, bringing quantum computing closer to practical applications.
@@
One particularly exciting and rapidly evolving domain within quantum computing is quantum machine learning (QML), which lies at the intersection of quantum computing and machine learning [9, 10, 11, 12, 13]. QML aims to leverage quantum algorithms and data structures to tackle complex problems and achieve breakthroughs that classical machine learning methods struggle to address [14, 15, 16, 17, 18, 19].
In this paper, we explore the Quantum Extreme Learning Machine (QELM) [20, 21, 22, 23, 24, 25, 26, 27], a model that utilizes quantum dynamics to perform complex data transformations. The QELM builds on the Extreme Learning Machine (ELM) [28, 29, 30, 31, 32, 33, 34] framework, a classical machine learning model designed to streamline the training process by limiting learning to the output layer. In the QELM, this approach is extended into the quantum realm, where a fixed-structure quantum reservoir processes input data and maps it into a high-dimensional space, yielding rich representations of the data.
Unlike classical Extreme Learning Machines, the QELM harnesses the vast state space of quantum systems, enabling it to handle complex datasets with computational power and efficiency that surpasses traditional methods [16]. Additionally, while classical ELMs are constrained by classical linear algebra, the QELM exploits the high dimensionality of quantum systems, making it highly effective for pattern recognition and classification tasks [9].
The input data is first encoded into quantum states using specific encoding schemes, which are then fed into a quantum reservoir. This reservoir (a system with fixed internal dynamics) transforms the input into higher-dimensional representations where a larger distinguishability is possible. Importantly, the quantum reservoir operates without requiring optimization or training of its internal parameters, allowing complex data transformations to occur naturally. Once processed, the data is passed to the output layer, where learning occurs. Only the weights of this layer are optimized, which significantly reduces training complexity and computational cost. One of the primary advantages of the QELM is its efficiency in training. Unlike many machine learning models that rely on backpropagation and iterative adjustment of all parameters, the QELM confines learning to the output layer. This approach greatly reduces computational overhead, resulting in faster training times and lower resource consumption.
Furthermore, the quantum reservoir’s ability to map input data into a vast Hilbert space, which grows exponentially with the number of quantum particles, allows the QELM to capture intricate relationships within the data, revealing patterns that classical methods might miss. This ability to handle complex, high-dimensional transformations is particularly valuable in domains where feature extraction and pattern recognition are key challenges.
QELMs fall within the broader framework of Quantum Reservoir Computing (QRC) [35, 25, 36], but operate without memory, i.e. the reservoir’s dynamics are reduced to a static, high-dimensional input transformation [37]. While classical reservoir computing has been extensively studied in the context of recurrent neural networks (RNNs) [38, 39, 40, 41, 42, 43, 44] and physical systems [45, 46, 47, 48], several advances have extended this framework to quantum systems [35, 49, 50]. In particular, spin chains [51, 52, 53, 54, 55, 56, 57, 58, 59, 60], fermionic and bosonic systems [61, 62, 63, 64, 65, 66, 67], quantum oscillators [68, 69], and Rydberg atoms[70, 71, 24] have been explored as quantum physical implementations. Furthermore, to enhance the memory of previous input injections, hybrid approaches and feedback mechanisms have been proposed [72, 73, 74, 75, 76, 77]. Since QELMs do not retain information from previous inputs, they are less suitable for tasks requiring temporal dependencies or sequence learning, such as time-series forecasting.
However, this memoryless nature simplifies the model’s structure, enabling it to focus on instantaneous data transformations, which is ideal for tasks where temporal dependencies are not essential. Given its properties, the QELM is particularly well-suited for classification, regression, and pattern recognition tasks in domains that do not rely on time-dependency. In fields such as image recognition, fraud detection, and static data analysis, the ability to efficiently map input data to high-dimensional spaces using quantum resources can lead to significant performance improvements [78, 9, 16, 79, 80]. Furthermore, the rapid training process of the QELM makes it an attractive option for applications where time or computational resources are limited.
@@
Throughout this paper, we study the architecture and performance of QELM for image classification tasks. Due to the limitations arising from the reduced number of qubits available today and the small decoherence times of current noisy quantum devices [6, 81], typical datasets, characterized by a large number of features, cannot be directly fed to the quantum reservoirs and must necessarily be compressed into a lower-dimensional representation. In the context of QELM, this is usually achieved by linear reduction algorithms, such as Principal Component Analysis (PCA).
In this work, we perform a step forward and exploit the more complex representation power of nonlinear neural network architectures such as the Auto-Encoders. In particular, we quantitatively show that the reservoir performances are substantially improved when the Auto-Encoder, rather than PCA, is used to compress the number of features.
@@
Furthermore, since it is necessary to convert classical information into quantum states that the quantum system can process, we explore various encoding methods, identifying those that work better. Finally, we compare the results obtained from different Hamiltonians that generate the natural dynamics of the reservoir, gaining some insights into the learning processes and performances of the various reservoirs.
@@
The paper is structured as follows: Section 2 provides a description of the model used. Sections 3 and 4 cover techniques for reducing data complexity and methods for encoding classical input information, respectively. In Section 5, we present different Hamiltonian models. Section 6 details the measurement process and the classical machine learning classifier. Finally, Section 7 discusses the model’s performance and results, followed by the conclusion.
@@
To demonstrate that our findings are not limited to a specific dataset and are generally applicable, we also test our approach on the more complex Fashion-MNIST dataset [83], for which the same general considerations hold.
Furthermore, the accuracies obtained in our study are comparable to or exceed those of other image classification methods based on quantum machine learning, such as Quantum Capsule Networks [13], Quantum Convolutional Neural Networks [84], the QRC proposed by [24], the Quantum Bayes Classifiers proposed in [85] or the approaches proposed in [86, 87].
@@ 2   Overview of Quantum Extreme Learning Machines
We will explore a QELM model that combines a classical Extreme Learning Machine framework, in which the training process is restricted to the output layer only, with a quantum reservoir. As said before, we feed the algorithm with an image dataset (MNIST) to implement an image classification task. A schematic representation of the model is shown in Fig.1 and its workflow can be summarized as follows:
  • Feature reduction: the original data is compressed into a lower dimensional latent representation to match the limited number of available qubits of the reservoir. This can be achieved using either linear Principal Component Analysis (PCA) or a nonlinear Auto-Encoder (AE) [88, 89, 90, 91].
@@
Figure 1: A schematic representation of the QELM. The workflow is as follows: feature reduction through PCA or AE; encoding of classical data into the quantum initial state; time evolution via the quantum layer; measurement of the evolved quantum state; classification with a classical One-layer Neural Network.
@@ 3   Techniques for reducing data complexity
In depth, we examine two distinct approaches for reducing the number of dimensions of the features: the Principal Component Analysis (PCA) and the Auto-Encoder (AE).
@@
In particular, we used two different algorithms of auto-encoders that we name $AE_{1}$ and $AE_{2}$. Both utilize in the encoder convolutional and max-pooling layers for feature extraction, while convolutional and up-sampling layers are used in the decoder for the reconstruction of the input data. A fully-connected layer maps the aforementioned features into a latent-space representation which is then fed to the quantum algorithm. The fully-connected layer employs a sigmoid activation function (the outputs of which are real values between 0 and 1) while the convolutional layers rely on the rectified linear unit. The two auto-encoders mainly differ in their expressivity potential, with $AE_{1}$ characterized by a number of trainable parameters larger than the one in $AE_{2}$. More details on the architectures of $AE_{1}$ and $AE_{2}$ can be found in Appendix A.
@@ 4   Encoding strategies of classical data into the QELM
|$\ket{\vec{x}}=\bigotimes_{i=1}^{\left\lceil M/2\right\rceil}\left(\cos(\frac{x_{2i-1}}{2})\ket{0}+e^{ix_{2i}}\sin(\frac{x_{2i-1}}{2})\ket{1}\right)$ (2)
where $x_{i}$ represents the elements of the feature vector $\vec{x}=[x_{1},...,x_{M}]^{T}\in\mathbb{R}^{M}$ and $N=\left\lceil M/2\right\rceil$ is the number of qubits. We chose to normalize each feature in the interval $[0,\pi]$, thus effectively covering, for each qubit, half of the corresponding Bloch sphere [22], with the angles $\theta$ and $\phi$ identified with pairs of features ($x_{2i-1}$, $x_{2i}$).
@@
A simpler case is represented by a single qubit encoding which maps each feature into one qubit ($M=N$), without exploiting the relative phase, namely with $\phi=0$. In this case, the initial state is prepared according to
|$\ket{\vec{x}}=\bigotimes_{i=1}^{M}\left(\cos(\frac{x_{i}}{2})\ket{0}+\sin(\frac{x_{i}}{2})\ket{1}\right).$ (3)
@@
|$\ket{\vec{x}}=\bigotimes_{i=1}^{\left\lceil M/2\right\rceil}\left(\sqrt{x_{2i-1}}\ket{0}+e^{ix_{2i}}\sqrt{1-x_{2i-1}}\ket{1}\right)$ (4)
@@
|$\ket{\vec{x}}=\bigotimes_{i=1}^{\left\lceil M/2\right\rceil}\frac{1}{\sqrt{x_{2i-1}^{2}+x_{2i}^{2}}}\left(x_{2i-1}\ket{0}+x_{2i}\ket{1}\right).$ (5)
@@ 5   Hamiltonian models describing the quantum layer
In this section, with the aim of understanding the features of the dynamics that help to improve the performances, we explore the influence of various Hamiltonians generating qualitatively different dynamics of the qubit register. Specifically, we analyze six distinct Hamiltonians, which govern the time evolution of the initial state containing the encoded features, on the classification accuracy. In particular, we consider the time-dependent Hamiltonian suggested in [22], alongside five time-independent operators. Among the latter, two have the qubit lying on a fully connected graph, with all-to-all interactions, while three others model qubits arranged in a linear configuration, with only nearest neighbor interactions.
@@
Here, $\sigma_{x}^{(i)}$, $\sigma_{z}^{(i)}$, and $\sigma_{z}^{(j)}$ are the Pauli operators acting on the site of the $i$,$j$-th qubit, while $J_{1}^{i,j}=\frac{J_{0}}{|i-j|^{\alpha}}$ represents the long-range interaction between the $i$-th and $j$-th qubits. As in [22], the scale of the interaction is set by $J_{0}=0.06$ and is damped for far-away qubits, with $\alpha=1.51$ controlling the rate of damping. $B_{1}=3.05$ and the time evolution is performed using a $\Delta t=50T$ ($T=2T_{1}=1$, setting the units for both time and inverse energy).
In the second case, we use the Hamiltonian $H_{2}$ below, featuring interacting qubits in a transverse field, [81]
@@
Here, the coefficients $J_{2}^{i,j}$ and $B_{2}^{(i)}$ are sampled from the Gaussian distributions $N(0.75,0.1)$ and $N(1,0.1)$, respectively. The time evolution is implemented with a time $\Delta t=20$. [81]
In the third case, we consider a Ising-like model, with only nearest neighbor interaction, including both longitudinal and transverse fields, as used in [23],
@@
where we set the parameters in the chaotic regime: $J_{3}=-1$, $B_{3x}=0.7$ and $B_{3z}=1.5$. For $H_{3}$ as well, the time evolution is performed using a $\Delta t=20$[23].
As a variation of the latter and as our fourth example, we employ the Heisenberg XXZ model, whose Hamiltonian we dub $H_{4}$:
$H_{4}=-\frac{1}{2}\sum_{i=1}^{N}\Big{(}$ $J_{4x}\sigma^{(i)}_{x}\sigma^{(i+1)}_{x}+J_{4y}\sigma^{(i)}_{y}\sigma^{(i+1)}_{y}$
$+J_{4z}\sigma^{(i)}_{z}\sigma^{(i+1)}_{z}+B_{4z}\sigma^{(i)}_{z}\Big{)}$ (10)
where we set the parameters $J_{4x}=J_{4y}=2$ and $J_{4z}=B_{4z}=0.54$ and the time evolution is performed using a $\Delta t=20$.
In Ref. [53], the performance of QRC has been related to the localization and ergodicity properties of the Hamiltonian eigenspectrum. With the aim of testing the role of locality/ergodicity (and also of integrability) on the algorithm performance, we further explore two other kind of Hamiltonians; namely: a fully integrable model and the Hamiltonian reported in Ref. [53], which, depending on the values of its parameters, can display different localization features.
Thus, we consider the integrable and non-ergodic XX model, whose Hamiltonian describes (in the fermion language) just hopping between nearest neighbours, $H_{5}$:
$H_{5}=\frac{1}{2}\sum_{i=1}^{N}(\sigma_{x}^{(i)}\sigma_{x}^{(i+1)}+\sigma_{y}^{(i)}\sigma_{y}^{(i+1)}).$ (11)
As our final example, we take the Hamiltonian $H_{6}$ reported in Ref. [53]:
$H_{6}=\sum_{i>j=1}^{N}J_{6}^{i,j}\sigma^{(i)}_{x}\sigma^{(j)}_{x}+\frac{1}{2}\sum_{i=1}^{N}(B_{6}+D_{i})\sigma^{(i)}_{z}$ (12)
where $J_{6}^{i,j}$ are randomly selected from a uniform distribution in the interval $[-0.5,0.5]$, $D_{i}$ is randomly drawn from the uniform distribution $[-W,W]$. In detail, we analyzed the reservoir dynamics generated by $H_{6}$ using three different set of parameters, corresponding to non-ergodic regions described in [53]: a first one, where localization occurs (region III in Ref. [53]) with $B_{6}=W=2*10^{-2}$, a second one, laying in between non-ergodic and ergodic regions, with $B_{6}=0.03$ and $W=1$ and a third set of parameters lying in the many-body-localized region (region I of Ref. [53]), with $B_{6}=0.03$ and $W=60$.
Moreover, to help the comparison we have also taken into account the Hamiltonian $H_{2}$ with $J_{2}=0$, as this choice represents a system where there are no interactions between the spins.
The evolution generated by all of these Hamiltonians, implemented using the QuTiP package [97, 98], is taken for a time $\Delta t=20$.
@@ 7.1   PCA vs Autoencoders
As naively expected, the simpler linear PCA shows a lower accuracy compared to the neural-network-based algorithms, $AE_{1}$ and $AE_{2}$, and $AE_{1}$ performs better than $AE_{2}$ due to its higher capacity. We notice that, after a fast initial growth, the accuracy starts to saturate around $N=9$ qubits to $\sim 98\%$ for the test set and slightly higher for the training set. As the number of qubits increases, the differences among the classification performances of the three algorithms get reduced. This behaviour can be ascribed to the corresponding increase in the dimension of the latent representation, combined with the relative simplicity of the MNIST dataset. As such, we expect the gap between the autoencoders and the PCA to persist even for large $N$ for more complex datasets. As an example, we considered the Fashion-MNIST dataset and confirmed the presence of the gap. The corresponding results are provided in Appendix B.
Figure 2: Training (left panel) and testing (right panel) accuracy as a function of the number of qubits for different feature reduction schemes: PCA (blue) and two autoencoders ($AE_{1}$ and $AE_{2}$, green and red curves respectively). The time evolution and the encoding have been performed with the Hamiltonian $H_{1}$ and the dense angle, respectively.